Monterrey Properties enters into a 4-year lease for an automobile to be used in operation. The agreement obligates the company to make lease payments of $10,000 at the end of every year. Its useful economic life would be 8 years. Assume that the company’s borrowing rate is 6%. The PV factor for 4 period, 6%, ordinary annuity is 3.4651.

What expenses will Monterrey record for thefirst year of the lease under IFRS?

A.

Depreciation expense for $10,000.

B.

Lease expense for $10,000.

C.

Depreciation expense for $8,663 and Interest expense for $10,000.

D.

Depreciation expense for $8,663 and Interest expense for $2,079.
Answer #1

Lease payable =10000*3.4651 = $34651

Interest expense = 34651*6% = $2079

Depreciation = 34651/4 = $8663

Option D. is correct answer.